OpenStudy (anonymous):

"Online purchasers of a computer game A also often bought two other games L and H. A recent study has shown that 16% of purchasers of A also bought L; 10% also bought H and that 10% of purchasers of one of H and L also bought the other. If a purchaser of A is found to have bought H; what is the probability that they also bought L?"

OpenStudy (anonymous):

This is all I have for now :( \[P(L|A) = 0.16\]\[P(H|A) = 0.10\]\[P(H|L) + P(L|H) = 0.10\]

OpenStudy (anonymous):

\[P(L|A,H) = \]

OpenStudy (anonymous):

\[P(L|A\cap H) = \]I'm a little stuck....

OpenStudy (anonymous):

give me a hint :(

OpenStudy (anonymous):

I should probably compute the probabilty that someone bought A first

OpenStudy (anonymous):

Ahh that hint means I don't have to :-P

OpenStudy (anonymous):

The problem is intended to be entirely about purchasers of a computer game A and what other games they might purchase. You can therefore treat purchasers of the game A as the universal set. The 16% and 10% for L and H respectively are intended as independent figures

OpenStudy (anonymous):

\[P(H) = 0.16\]\[P(L) = 0.10\]\[P(L|H) + P(H|L) = 0.10\] that looks better

OpenStudy (anonymous):

\[P(L|H) = P(L)\]

OpenStudy (anonymous):

!!! :(

OpenStudy (anonymous):

Let's check that answer, and run a Python simulation :-D

OpenStudy (anonymous):

or is it wrong?

OpenStudy (anonymous):

Try posting it here; they love problems like these over there http://math.stackexchange.com/

OpenStudy (anonymous):

I will go shower and come back

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

I wish JamesJ would help.... I'm extremely sure I'm looking at this problem the wrong way, and that it is actually an elementary application of probability laws :-P

OpenStudy (anonymous):

Alright I will try the python solution.

OpenStudy (jamesj):

Ok, so what you actually have is P(L|A)=0.16 P(H|A)=0.10 P(H|L)=0.10 P(L|H)=0.10 and you want to know P(L|H,A)

OpenStudy (mr.math):

Is P(L|H,A)=P(L|H and A)?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

H intersect A :-D

OpenStudy (jamesj):

this is tricky. Here's what I've got.

OpenStudy (jamesj):

My strategy is to find P(L|HA) as part of a complete probability calculation P(L|A) = P(L|HA)P(H|A) + P(L|-HA)P(-H|A) Now we know P(L|A), P(H|A) and P(-H|A). We want P(L|HA) and we could get it except for this problem term of P(L|-HA). Following so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

on the python simulation I wrote I ran into 0.01 as the number of people who first bought H and then bought L O.o

OpenStudy (jamesj):

So let's expand P(L|-HA) P(L|-HA) = P(-H|LA)P(L|A)/P(-H|A) = ( 1 - P(H|LA) ) P(L|A) / (1 - P(H|A) Now we know every term except P(H|LA)

OpenStudy (jamesj):

so one more crank of the Bayes Rule handle: P(H|LA) = P(L|HA)P(H|A)/P(L|A) Now P(L|HA) is the quantity we want to find. Put all of this back together and it should be the only variable in the complete probability equation. Solve for it.

OpenStudy (anonymous):

P(L|A)=0.16 P(H|A)=0.10 P(H|L)=0.10 P(L|H)=0.10 punching those in the calculator :-D

OpenStudy (anonymous):

You guys are trying to make my brain explode by reading that stuff... There should be a warning label

OpenStudy (anonymous):

I got 8/5 O.o

OpenStudy (anonymous):

Here is my calculation : \[0.16 = x*0.10 + \frac{0.16-x*0.10}{1-0.10}\]

OpenStudy (anonymous):

This is the python script I wrote.... I think it's also wrong http://ideone.com/Dr4xd

OpenStudy (anonymous):

in the script, I had 1000000 purchasers of game A.

OpenStudy (jamesj):

I see what happens. Actually all the algebra collapses to a the trivial equation, 0 = 0.

OpenStudy (anonymous):

lol

OpenStudy (jamesj):

Which is perhaps to be expected, applying Bayes Rule twice to the same term. We need a new strategy

OpenStudy (anonymous):

I think the A part is a red herring. What if we treat the universal set just as 'people' and then draw subsets of buyers of H, buyers of L etc. from that

OpenStudy (anonymous):

or am I mistaken?

OpenStudy (jamesj):

No, buying A does give you information on buying L or H. In the Bayes network of L, H and A, A is at the top of triangle with arrows pointing down to L and H, and L and H have arrows going each way from one to the other

OpenStudy (anonymous):

right

OpenStudy (anonymous):

is that the right model of the problem? That deciding to buy a game is influenced by a previous decision to buy a different game?

OpenStudy (jamesj):

It's actually not helpful to think of this in a Bayes network; I was just using that as an example of the dependencies here. All I want to say in particular is it is NOT true that P(L|HA) = P(L|H)

OpenStudy (anonymous):

right

OpenStudy (jamesj):

I don't know. I'm at a dead end right now.

OpenStudy (jamesj):

Let me think about it and get back to you. It's worth asking Zarkon as well, as he works in probability and statistics, while I'm just a visitor with a good travel guide.

OpenStudy (anonymous):

I'll see if I can post this one at the stackexchange.

OpenStudy (jamesj):

yes, and let me know if you get a good answer. Where does this problem come from btw?

OpenStudy (anonymous):

Tomas.A just asked me in one of those 'Fan Testimonial' things.

OpenStudy (anonymous):

lol

OpenStudy (jamesj):

Thomas just sent me this link, that unravels the riddle of this problem: http://pastebin.com/s4Erdcfi

OpenStudy (zarkon):

ah

OpenStudy (jamesj):

@Thomas, post your question here. But look at this, quoting from the teacher's response: "The problem in Q1 is intended to be entirely about purchasers of a computer game A and what other games they might purchase. You can therefore treat purchasers of the game A as the universal set. The 16% and 10% for L and H respectively are intended as independent figures" However again a lot of students complained that it's unclear and he said ( worth mentioning that he write it only 15mins before deadline for all assignments): "There are 2 sets involved A: those who bought one (at least) of the games B: those who bought both B comprises 10% OF A" Hence in fact P(H|LA) = P(H|L) , because A is the universal set here, or formally P(A) = 1 = P(H) , because in fact H and L are independent = 0.10 ********** I put this down as a poorly phrased question.

OpenStudy (anonymous):

so the answer to the question is 0.10 :-D phew!